# SCoT: Straight Consistent Trajectory for Pre-Trained Diffusion Model Distillations

### v3

@@ Abstract
Pre-trained diffusion models are commonly used to generate clean data (e.g., images) from random noises, effectively forming pairs of noises and corresponding clean images. Distillation on these pre-trained models can be viewed as the process of constructing advanced trajectories within the pair to accelerate sampling. For instance, consistency model distillation develops consistent projection functions to regulate trajectories, although sampling efficiency remains a concern. Rectified flow method enforces straight trajectories to enable faster sampling, yet relies on numerical ODE solvers, which may introduce approximation errors. In this work, we bridge the gap between the consistency model and the rectified flow method by proposing a Straight Consistent Trajectory (SCoT) model. SCoT enjoys the benefits of both approaches for fast sampling, producing trajectories with consistent and straight properties simultaneously. These dual properties are strategically balanced by targeting two critical objectives: (1) regulating the gradient of SCoT’s mapping to a constant, (2) ensuring trajectory consistency. Extensive experimental results demonstrate the effectiveness and efficiency of SCoT.
@@ 1 Introduction
Pre-trained diffusion models Ho et al. (2020); Song et al. (2021b); Rombach et al. (2022); Poole et al. (2023); Esser et al. (2024) have demonstrated impressive performance in real-world tasks such as high-quality image synthesis and image editing. However, such practical models usually require extensive computational resources to train, as well as a large number of model evaluations to generate high-quality samples (e.g. images). Leveraging these pre-trained teacher models to generate pairs of random noises and their corresponding clean images, a popular choice for low-cost training and fast sampling is to train a student distillation model with advanced “trajectories” within the pair Salimans and Ho (2022); Wang et al. (2023, 2024); Yin et al. (2024b); Luo et al. (2024); Xie et al. (2024); Nguyen and Tran (2024); Yin et al. (2024a); Sauer et al. (2024); Xu et al. (2024); Zhu et al. (2025); Frans et al. (2025a); Sauer et al. (2025).
@@
In order to address these issues and enjoy the benefits of both approaches, we bridge the gap between consistency models and rectified flows by proposing the Straight Consistent Trajectory (SCoT) model in this paper. SCoT aims to produce trajectories that are consistent, which is a valid condition for trajectories, and are straight, which simplifies the point changes in the trajectory. In detail, SCoT regulates the trajectory through two aspects: (1) trajectory straightness by optimizing velocities towards the amount of point changes in the pre-trained model, (2) trajectory consistency by ensuring points from different time steps are projected to converge at the same value at future time steps.
In summary, the proposed SCoT is the first to produce consistent and straight trajectories between random noise and clean images. Consistency guarantees valid trajectories, while straightness facilitates the approximation of the trajectory projection function and faster sample generation. With this design, SCoT achieves state-of-the-art results, enabling compact models to generate high-quality data in $N$-steps, or even a single step.  Figure˜1 highlights the key differences among consistency distillation, rectified flow distillation, ShortCut model Frans et al. (2025b), and the proposed SCoT.
@@ 2 A trajectory perspective on Pre-trained diffusion model distillations
Consistency distillation Consistency model (CM) Song et al. (2023) can be regarded as one trajectory distillation method. CM studies a consistent projection function $f_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t)$ that maps any noisy image ${\boldsymbol{\mathbf{x}}}_{t}$ to the clean image ${\boldsymbol{\mathbf{x}}}_{0}$: $f_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t)={\boldsymbol{\mathbf{x}}}_{0},\forall t\in[0,1]$. Its distillation objective function can be written as a weighted distance metric ${\boldsymbol{\mathbf{D}}}(\cdot,\cdot)$ between the mapped clean images from two adjacent points $\mathcal{L}_{\text{CM}}=\mathbb{E}_{t}\left[\omega(t){\boldsymbol{\mathbf{D}}}\left({\boldsymbol{\mathbf{f}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t+\Delta t},t+\Delta t),{\boldsymbol{\mathbf{f}}}_{{\boldsymbol{\mathbf{\phi}}}^{-}}(\widehat{{\boldsymbol{\mathbf{x}}}}^{{\boldsymbol{\mathbf{\theta}}}}_{t},t)\right)\right]$, where $\boldsymbol{\mathbf{\phi}}$$^{-}$ is the exponential moving average of the past values $\boldsymbol{\mathbf{\phi}}$}, and $\widehat{{\boldsymbol{\mathbf{x}}}}^{{\boldsymbol{\mathbf{\theta}}}}_{t}$ is obtained from pre-trained model as $\widehat{{\boldsymbol{\mathbf{x}}}}^{{\boldsymbol{\mathbf{\theta}}}}_{t}={\boldsymbol{\mathbf{x}}}_{t+\Delta t}-t\Delta t\nabla_{{\boldsymbol{\mathbf{x}}}_{t+\Delta t}}\log p_{t+\Delta t}({\boldsymbol{\mathbf{x}}}_{t+\Delta t})$. CM obtains a “valid” trajectory by mapping different points to the same clean image.
@@
3 Straight Consistent Trajectory (SCoT) model
Given a pre-trained noise-image pair $({\boldsymbol{\mathbf{x}}}_{1},\widehat{{\boldsymbol{\mathbf{x}}}}_{0})$, the proposed Straight Consistent Trajectory (SCoT) model aims to learn a trajectory-based projection function $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ that produces straight and consistent trajectories within the pair. Similar to the CTM, the projection function $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ takes the current time step $t$, the values ${\boldsymbol{\mathbf{x}}}_{t}$ at step $t$, as well as the future time step $s$ as inputs, and outputs the values at time step $s$ as $\widehat{{\boldsymbol{\mathbf{x}}}}_{s}:=G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$.
@@ 3.1 Main Components of SCoT
Let ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$ denote the velocity of the student trajectory at time step $s$ and let ${\boldsymbol{\mathbf{x}}}_{t}$ denote the value at initial step $t$, $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ is equivalent to the solution of the ODE ${\mathrm{d}\widehat{{\boldsymbol{\mathbf{x}}}}_{s}}/{\mathrm{d}s}={\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}(\widehat{{\boldsymbol{\mathbf{x}}}}_{s},s),\forall s\in[t,1]$. That is, $\widehat{{\boldsymbol{\mathbf{x}}}}_{s}=G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)={\boldsymbol{\mathbf{x}}}_{t}+\int_{t}^{s}{\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{r},r)\mathrm{d}r$. Since ${\partial G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)}/{\partial s}={\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$, Equation˜2 can be alternatively understood as enforcing the student trajectory’s velocity ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$ close to the magnitude of changes in the pre-trained teacher model.
@@
Figure 2: From two different points ${\boldsymbol{\mathbf{x}}}_{t_{1}},{\boldsymbol{\mathbf{x}}}_{t_{2}}$, SCoT maps into the same point $\widehat{{\boldsymbol{\mathbf{x}}}}_{s}$ with the same velocity ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}(\widehat{{\boldsymbol{\mathbf{x}}}}_{s},s)$ at time step $s$.
@@
where $\lambda_{\text{vel}},\lambda_{\text{con}}$ denote the weighting factors for $\mathcal{L}_{\text{velocity}},\mathcal{L}_{\text{consistency}}$.
@@ 3.3 Connections to Other Distillation Methods
The Shortcut model Frans et al. (2025a) shares the same targets of trajectory consistency and straightness. While emphasizing learning trajectory velocities instead of the full trajectory, the Shortcut model still relies on an ODE solver for image sampling. Also, its usage of a one-step Euler solver to enforce trajectory consistency might not be optimal.
@@ 4.1 Experimental Setup
Teacher-Student Distillation Setup. To simulate the generation trajectory from $t=0$ to $t=1$, the pre-trained models are used to generate paired data samples $\langle{\boldsymbol{\mathbf{x}}}_{0},{\boldsymbol{\mathbf{x}}}_{1}\rangle$. Furthermore, to provide the intermediate states required for computing the consistency term in  Equation˜3, we distill knowledge from a teacher model $\boldsymbol{\mathbf{\theta}}$} into a student model $\boldsymbol{\mathbf{\phi}}$}. Following established practices in CM Song et al. (2023), SlimFlow Zhu et al. (2025), and CTM Kim et al. (2024), we evaluate SCoT on both CIFAR-10 and ImageNet, using pre-trained diffusion checkpoints from EDM [^1] https://github.com/NVlabs/edm (for CIFAR-10) and CM [^2] https://github.com/openai/consistency_models (for ImageNet) as teacher models. For the student model, we adopt EDM’s DDPM++ implementation on CIFAR-10, and the Ablated Diffusion Model (ADM) architecture from Dhariwal and Nichol (2021) for ImageNet. To support the additional time conditioning $s$ in  Equation˜4, we incorporate several architectural modifications inspired by CTM. Specifically, we extend the temporal embedding by adding auxiliary $s$-embedding information to the original $t$-embedding. We also adopt other training heuristics from CTM (see Table˜8), including: (1) using a larger value of $\mu$ in the stop-gradient EMA to slow the teacher update rate and improve training stability; (2) setting the student EMA rate to 0.999 to smooth parameter updates over time and reduce training noise; and (3) reusing skip connections from the pre-trained diffusion model to facilitate gradient flow and preserve hierarchical features.
@@
Model Architectures. To support the additional time conditioning variable $s$ in our generator $g_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}},t,s)$, we incorporate a conditional embedding module into the baseline model architectures, following prior design choices in CM Song et al. (2023), SlimFlow Zhu et al. (2025), and CTM Kim et al. (2024). For CIFAR-10, we adopt the DDPM++ implementation from EDM Karras et al. (2022), extending its time embedding module to additionally encode $s$ using a structure symmetric to the original $t$ embedding, and applying identical normalization strategies. For ImageNet, we use the Ablated Diffusion Model (ADMDhariwal and Nichol (2021), where $s$ is embedded jointly with the class conditioning variable $c$, requiring no architectural change. A detailed comparison of DDPM++ and ADM in terms of their ResNet backbones, attention configurations, and noise conditioning strategies is presented in Table˜9. To improve the attention module, we address compatibility issues in QKVFlashAttention by modifying the dimension operations in QKVAttentionLegacy to match expected checkpoint formats. Additionally, we integrate xformersScaledDotProduct attention as an alternative backend. These improvements ensure correct weight loading while enhancing efficiency and flexibility of the attention mechanism.
@@
Table 2: Comparison of NN-step generation performance across diffusion models on CIFAR-10 at comparable model scales. We report sample quality metrics—FID ($\downarrow$)-for $N\in\{1,2,10,20,50\}$. Entries highlighted in bold denote our proposed method. CTM indicates the inclusion of GAN loss. Baseline results are sourced from Song et al. (2023); Zhu et al. (2025); Frans et al. (2025b). “–” indicates that the result is not available.
Training and Sampling Hyperparameters. To ensure stable convergence and fair comparison during the training of SCoT, we adopt a higher learning rate for smaller datasets (e.g., CIFAR-10) and scale the batch size appropriately for larger datasets (e.g., ImageNet). We use the Adam optimizer for both settings, and enable mixed-precision training (FP16) to improve memory efficiency and computational speed. For CIFAR-10, we train the model for 130k iterations with a batch size of 512, while for ImageNet we use a larger batch size of 2048 and train for 40k iterations to accommodate the increased data complexity and training cost. We adopt an exponential moving average (EMA) of the student model with a decay rate of 0.9999 and employ a stop-gradient variant with EMA coefficient $\mu=0.9999$ to stabilize the training objective. A complete list of hyperparameter settings is provided in Table˜8.
Datasets. For evaluation, we adopt two large-scale real-world datasets with different image resolutions: CIFAR-10 ($32\times 32$) and ImageNet ($64\times 64$), following standard protocol. Dataset statistics, including resolution, total size, and sample count, are summarized in Table˜10.
Evaluation Metrics. We assess SCoT on unconditioned image generation using standard evaluation metrics, including Fréchet Inception Distance (FID) Heusel et al. (2017), Negative Log-Likelihood (NLL), Inception Score (IS) Salimans et al. (2016), and Recall (Rec.) Sajjadi et al. (2018). To evaluate model efficiency, we also report parameter count, floating-point operations (FLOPs), and multiply–accumulate operations (MACs). See Section˜C.2 for further details.
Time Efficiency. We measure the training throughput of the trajectory generator $g_{{\boldsymbol{\mathbf{\phi}}}}(\cdot)$ under different combinations of loss functions to evaluate time efficiency. Throughput results, reported in images per second per GPU, are summarized in  Table˜11.

### v4

# SCoT: Unifying Consistency Models and Rectified Flows via Straight-Consistent Trajectories

@@ Abstract
Pre-trained diffusion models are commonly used to generate clean data (e.g., images) from random noises, effectively forming pairs of noises and corresponding clean images. Distillation on these pre-trained models can be viewed as the process of constructing advanced trajectories within the pair to accelerate sampling. For instance, consistency model distillation develops consistent projection functions to regulate trajectories, although sampling efficiency remains a concern. Rectified flow method enforces straight trajectories to enable faster sampling, yet relies on numerical ODE solvers, which may introduce approximation errors. In this work, we bridge the gap between the consistency model and the rectified flow method by proposing a Straight-Consistent Trajectories (SCoT) model. SCoT enjoys the benefits of both approaches for fast sampling, producing trajectories with consistent and straight properties simultaneously. These dual properties are strategically balanced by targeting two critical objectives: (1) regulating the gradient of SCoT’s mapping function to a constant and (2) ensuring trajectory consistency. Extensive experimental results demonstrate the effectiveness and efficiency of SCoT.
@@ 1 Introduction
Pre-trained diffusion models Ho et al. (2020); Song et al. (2021b); Rombach et al. (2022); Poole et al. (2023); Esser et al. (2024) have demonstrated impressive performance in real-world tasks such as high-quality image synthesis and image editing. However, such practical models usually require extensive computational resources to train, as well as a large number of model evaluations to generate high-quality samples (e.g. images). Using these pre-trained teacher models to generate pairs of random noises and their corresponding clean images, a popular choice for low-cost training and fast sampling is to train a student distillation model with advanced “trajectories” within the pair Salimans and Ho (2022); Wang et al. (2023, 2024); Yin et al. (2024b); Luo et al. (2024); Xie et al. (2024); Nguyen and Tran (2024); Yin et al. (2024a); Sauer et al. (2024); Xu et al. (2024); Zhu et al. (2025); Frans et al. (2025a); Sauer et al. (2025).
@@
In order to address these issues and enjoy the benefits of both approaches, we bridge the gap between consistency models and rectified flows by proposing the Straight-Consistent Trajectories (SCoT) model in this paper. SCoT aims to produce trajectories that are consistent, which is a valid condition for trajectories, and are straight, which simplifies the point changes in the trajectory. In detail, SCoT regulates the trajectory through two aspects: (1) trajectory straightness by optimizing velocities towards the amount of point changes in the pre-trained model, (2) trajectory consistency by ensuring points from different time steps are projected to converge at the same value at future time steps.
In summary, the proposed SCoT is the first to produce consistent and straight trajectories between random noise and clean images, which unifies the consistency models and rectified flows. Consistency guarantees valid trajectories, while straightness facilitates the approximation of the trajectory projection function and faster sample generation. With this design, SCoT achieves state-of-the-art results, enabling compact models to generate high-quality data in $N$-steps, or even a single step.  Figure˜1 highlights the key differences among consistency distillation, rectified flow distillation, ShortCut model Frans et al. (2025b), and the proposed SCoT.
@@ 2 A trajectory perspective on Pre-trained diffusion model distillations
Consistency distillation Consistency model (CM) Song et al. (2023) can be regarded as one trajectory distillation method. CM studies a consistent projection function $f_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t)$ that maps any noisy image ${\boldsymbol{\mathbf{x}}}_{t}$ to the clean image ${\boldsymbol{\mathbf{x}}}_{0}$: $f_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t)={\boldsymbol{\mathbf{x}}}_{0},\forall t\in[0,1]$. Its distillation objective function can be written as a weighted distance metric ${\boldsymbol{\mathbf{D}}}(\cdot,\cdot)$ between the mapped clean images from two adjacent points $\mathcal{L}_{\text{CM}}=\mathbb{E}_{t}\left[\omega(t){\boldsymbol{\mathbf{D}}}\left({\boldsymbol{\mathbf{f}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t+\Delta t},t+\Delta t),{\boldsymbol{\mathbf{f}}}_{{\boldsymbol{\mathbf{\phi}}}^{-}}(\widehat{{\boldsymbol{\mathbf{x}}}}^{{\boldsymbol{\mathbf{\theta}}}}_{t},t)\right)\right]$, where $\boldsymbol{\mathbf{\phi}}$$^{-}$ is the exponential moving average of the past values $\boldsymbol{\mathbf{\phi}}$}, and $\widehat{{\boldsymbol{\mathbf{x}}}}^{{\boldsymbol{\mathbf{\theta}}}}_{t}$ is obtained from the pre-trained model as $\widehat{{\boldsymbol{\mathbf{x}}}}^{{\boldsymbol{\mathbf{\theta}}}}_{t}={\boldsymbol{\mathbf{x}}}_{t+\Delta t}-t\Delta t\nabla_{{\boldsymbol{\mathbf{x}}}_{t+\Delta t}}\log p_{t+\Delta t}({\boldsymbol{\mathbf{x}}}_{t+\Delta t})$. CM obtains a “valid” trajectory by mapping different points to the same clean image.
@@
Concurrent works MeanFlow Geng et al. (2025) is a concurrent work to SCoT that also unifies consistency models and rectified flows via velocity integration. Its key distinction lies in introducing a correction term derived from the gradient with respect to the starting step $t$, unlike SCoT’s use of the gradient at the terminating step $s$. Although rectified flow is regarded as a special case of MeanFlow under an infinitesimal integration range, it is not clear if MeanFlow itself can learn straight trajectories in practice. Another concurrent work FlowMap Sabour et al. (2025); Boffi et al. (2025) also shares the same target. By using the CTM as the backbone, it is unclear if such a setting can achieve both straight and consistent trajectories in practice.
3 Straight-Consistent Trajectories (SCoT) model
Given a pre-trained noise-image pair $({\boldsymbol{\mathbf{x}}}_{1},\widehat{{\boldsymbol{\mathbf{x}}}}_{0})$, the proposed Straight-Consistent Trajectories (SCoT) model aims to learn a trajectory-based projection function $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ that produces straight and consistent trajectories within the pair. Similar to the CTM, the projection function $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ takes the current time step $t$, the values ${\boldsymbol{\mathbf{x}}}_{t}$ at step $t$, as well as the future time step $s$ as inputs, and outputs the values at time step $s$ as $\widehat{{\boldsymbol{\mathbf{x}}}}_{s}:=G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$.
@@ 3.1 Main Components of SCoT
Let ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$ denote the velocity of the student trajectory at time step $s$ and let ${\boldsymbol{\mathbf{x}}}_{t}$ denote the value at initial step $t$, $G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)$ is equivalent to the solution of the ODE ${\mathrm{d}\widehat{{\boldsymbol{\mathbf{x}}}}_{s}}/{\mathrm{d}s}={\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}(\widehat{{\boldsymbol{\mathbf{x}}}}_{s},s),\forall s\in[t,1]$. That is:
$\boxed{\widehat{{\boldsymbol{\mathbf{x}}}}_{s}=G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)={\boldsymbol{\mathbf{x}}}_{t}+\int_{t}^{s}{\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{r},r)\mathrm{d}r\Rightarrow\frac{\partial G_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{t},t,s)}{\partial s}={\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)}$ (3)
Equation˜2 can be alternatively understood as enforcing the student trajectory’s velocity ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}}_{s},s)$ close to the magnitude of changes in the pre-trained teacher model.
@@
Figure 2: From two different points ${\boldsymbol{\mathbf{x}}}_{t_{1}},{\boldsymbol{\mathbf{x}}}_{t_{2}}$, SCoT maps to the same point $\widehat{{\boldsymbol{\mathbf{x}}}}_{s}$. The velocity ${\boldsymbol{\mathbf{\mu}}}_{{\boldsymbol{\mathbf{\phi}}}}(\widehat{{\boldsymbol{\mathbf{x}}}}_{s},s)$ at time step $s$ is independent from previous time steps $t_{1},t_{2}$.
@@
where $\lambda_{\text{vel}},\lambda_{\text{con}}$ denote the weighting factors for $\mathcal{L}_{\text{velocity}},\mathcal{L}_{\text{consistency}}$. Each individual component addresses the straightness and consistency of trajectories, respectively. In this way, the proposed SCoT successfully unifies the consistency models and rectified flows.
The two loss terms, $\mathcal{L}_{\text{velocity}}$ and $\mathcal{L}_{\text{consistency}}$, have different dependencies on the pre-trained model. Specifically, $\mathcal{L}_{\text{velocity}}$ operates on a data-noise pair $(\widehat{{\boldsymbol{\mathbf{x}}}}_{0},{\boldsymbol{\mathbf{x}}}_{1})$, while $\mathcal{L}_{\text{consistency}}$ requires the model’s parameters. To cope with their requirements, we propose generating the data point $\widehat{{\boldsymbol{\mathbf{x}}}}_{0}$ dynamically for each training step using the pre-trained model $\widehat{{\boldsymbol{\mathbf{x}}}}_{0}={\boldsymbol{\mathbf{x}}}_{1}+\int_{1}^{0}{\boldsymbol{\mathbf{v}}}_{{\boldsymbol{\mathbf{\theta}}}}({\boldsymbol{\mathbf{x}}}_{t},t)dt$. This avoids the need for pre-computed pairs. An additional potential improvement involves using a shared random noise vector ${\boldsymbol{\mathbf{x}}}_{1}$ for both losses, which may enhance training consistency when sampling intermediate points. We leave a detailed study of this method to future work.
@@ 3.3 Connections to Other Distillation Methods
The Shortcut model Frans et al. (2025a) shares the same targets of trajectory consistency and straightness. While emphasizing learning trajectory velocities instead of the full trajectory, the Shortcut model still relies on an ODE solver for image sampling. Also, its usage of a one-step Euler solver to enforce trajectory consistency might not be optimal. Other concurrent works such as MeanFlow Geng et al. (2025) and FlowMap Boffi et al. (2025); Sabour et al. (2025) have been discussed in Section 2.
@@ 4.1 Experimental Setup
To simulate the generation trajectory from $t=0$ to $t=1$, the pre-trained models are used to generate paired data samples $\langle{\boldsymbol{\mathbf{x}}}_{0},{\boldsymbol{\mathbf{x}}}_{1}\rangle$. Furthermore, to provide the intermediate states required for computing the consistency term in  Equation˜4, we distill knowledge from a teacher model $\boldsymbol{\mathbf{\theta}}$} into a student model $\boldsymbol{\mathbf{\phi}}$}. Following established practices in CM Song et al. (2023), SlimFlow Zhu et al. (2025), and CTM Kim et al. (2024), we evaluate SCoT on both CIFAR-10 and ImageNet, using pre-trained diffusion checkpoints from EDM [^1] https://github.com/NVlabs/edm (for CIFAR-10) and CM [^2] https://github.com/openai/consistency_models (for ImageNet) as teacher models. For the student model, we adopt EDM’s DDPM++ implementation on CIFAR-10, and the Ablated Diffusion Model (ADM) architecture from Dhariwal and Nichol (2021) for ImageNet. To support the additional time conditioning $s$ in  Equation˜5, we incorporate several architectural modifications inspired by CTM. Specifically, we extend the temporal embedding by adding auxiliary $s$-embedding information to the original $t$-embedding. We also adopt other training heuristics from CTM (see Table˜8), including: (1) using a larger value of $\mu$ in the stop-gradient EMA to slow the teacher update rate and improve training stability; (2) setting the student EMA rate to 0.999 to smooth parameter updates over time and reduce training noise; and (3) reusing skip connections from the pre-trained diffusion model to facilitate gradient flow and preserve hierarchical features.
Model Architectures.
@@
To support the additional time conditioning variable $s$ in our generator $g_{{\boldsymbol{\mathbf{\phi}}}}({\boldsymbol{\mathbf{x}}},t,s)$, we incorporate a conditional embedding module into the baseline model architectures, following prior design choices in CM Song et al. (2023), SlimFlow Zhu et al. (2025), and CTM Kim et al. (2024). For CIFAR-10, we adopt the DDPM++ implementation from EDM Karras et al. (2022), extending its time embedding module to additionally encode $s$ using a structure symmetric to the original $t$ embedding, and applying identical normalization strategies. For ImageNet, we use the Ablated Diffusion Model (ADMDhariwal and Nichol (2021), where $s$ is embedded jointly with the class conditioning variable $c$, requiring no architectural change. A detailed comparison of DDPM++ and ADM in terms of their ResNet backbones, attention configurations, and noise conditioning strategies is presented in Table˜9. To improve the attention module, we address compatibility issues in QKVFlashAttention by modifying the dimension operations in QKVAttentionLegacy to match expected checkpoint formats. Additionally, we integrate xformersScaledDotProduct attention as an alternative backend. These improvements ensure correct weight loading while enhancing efficiency and flexibility of the attention mechanism.
@@
Table 2: Comparison of NN-step (NFE) generation performance across diffusion models on CIFAR-10 at comparable model scales. We report sample quality metrics—FID ($\downarrow$)-for $N\in\{1,2,10,20,50\}$. Entries highlighted in bold denote our proposed method. CTM indicates the inclusion of GAN loss. Baseline results are sourced from Song et al. (2023); Zhu et al. (2025); Frans et al. (2025b). “–” indicates that the result is not available.
Training and Sampling Hyperparameters.
To ensure stable convergence and fair comparison during the training of SCoT, we adopt a higher learning rate for smaller datasets (e.g., CIFAR-10) and scale the batch size appropriately for larger datasets (e.g., ImageNet). We use the Adam optimizer for both settings, and enable mixed-precision training (FP16) to improve memory efficiency and computational speed. For CIFAR-10, we train the model for 130k iterations with a batch size of 512, while for ImageNet we use a larger batch size of 2048 and train for 40k iterations to accommodate the increased data complexity and training cost. We adopt an exponential moving average (EMA) of the student model with a decay rate of 0.9999 and employ a stop-gradient variant with EMA coefficient $\mu=0.9999$ to stabilize the training objective. A complete list of hyperparameter settings is provided in Table˜8.
Datasets.
For evaluation, we adopt two large-scale real-world datasets with different image resolutions: CIFAR-10 ($32\times 32$) and ImageNet ($64\times 64$), following standard protocol. Dataset statistics, including resolution, total size, and sample count, are summarized in Table˜10.
Evaluation Metrics.
We assess SCoT on unconditioned image generation using standard evaluation metrics, including Fréchet Inception Distance (FID) Heusel et al. (2017), Negative Log-Likelihood (NLL), Inception Score (IS) Salimans et al. (2016), and Recall (Rec.) Sajjadi et al. (2018). To evaluate model efficiency, we also report parameter count, floating-point operations (FLOPs), and multiply–accumulate operations (MACs). See Section˜C.2 for further details.
Time Efficiency.
We measure the training throughput of the trajectory generator $g_{{\boldsymbol{\mathbf{\phi}}}}(\cdot)$ under different combinations of loss functions to evaluate time efficiency. Throughput results, reported in images per second per GPU, are summarized in  Table˜11.